### v2

# More on Round-Robin Tournament Models with a Unique Maximum Score

Abstract
In this note, we extend a recent result showing the uniqueness of the maximum score in a classical round-robin tournament to round-robin tournament models with equally strong players, where the scores are valued on a lattice.
Keywords: Complete graph, large deviations, local limit theorem, maximal score, negative dependence
MSC2020: 05C20, 05C07, 60F10, 60F15
1 Introduction and Problem Statement
In a round-robin tournament, each of $n$ players competes against each of the other n1n-1 players. When player $i$ plays against player $j$, player $i$’s reward is a random variable $X_{ij}$. Let
$s_{i}(n)=\sum_{j=1,j\neq i}^{n}{X_{ij}}$ (1)
denote the score of player $i,1\leq i\leq n$, after playing against all the other n1n-1 players. We assume that all $n\choose 2$} pairs $\left(X_{ij},X_{ji}\right),1\leq i<j\leq n$ are independent. We refer to $\left(s_{1}(n),s_{2}(n),\ldots,s_{n}(n)\right)$ as the score sequence of the tournament.
Tournaments provide a model for a statistical inference called paired comparisons. Measuring players’ strength in chess tournaments by modeling paired comparisons of strength has a long history, first appearing in Zermelo (1929). An excellent introduction to the combinatorial and some probabilistic aspects (with citations to earlier works) is given in Moon (1968). The statistical aspects are presented in David (1988) (also with citations to earlier works). Additional material on round-robin tournaments and related issues can be found in recent papers and the references therein (Ross, 2022; Malinovsky and Moon, 2022).
In a classic round-robin tournament (Model $M_{1}$) for $i\neq j$, $X_{ij}+X_{ji}=1$, $X_{ij}\in\left\{0,1\right\}$, and $\mathsf{P}(X_{ij}=1)=\mathsf{P}(X_{ij}=0)=1/2$, i.e., each of $n$ players wins or loses a game against each of the other n1n-1 with equal probabilities. The tournament can be represented by complete oriented graphs in which the vertices represent the players, and each pair of distinct vertices $i$ and $j$ is joined by an edge oriented from $i$ to $j$ or from $j$ to $i$ according to whether, in the case of Model $M_{1}$, $X_{ij}=1$ or $X_{ji}=1$. Let $r_{n}(M_{1})$ denote the probability that the tournament with $n$ labeled vertices has a unique vertex with maximum score under model $M_{1}$. Epstein (1967) stated, without proof, that in a classical round-robin tournament (Model $M_{1}$ ), the probability that there is a unique vertex with the maximum score tends to $1$ as $n$ tends to infinity, i.e.,
$\lim_{n\rightarrow\infty}r_{n}(M_{1})=1.$} (2)
For values of $n$ that are very small, it is possible to compute $r_{n}(M_{1})$ from the distribution of the nondecreasing score sequences. However, even for small $n$, this becomes computationally impossible due to the fact that the size of the sample space is $2^{n\choose 2}$. Indeed, Epstein (1967) employed the distribution of the nondecreasing score vector up to $n=8$, as provided by David (1959), to calculate the values of $r_{1}(M_{1}),\ldots,r_{8}(M_{1})$. It is interesting to note that MacMahon (1923a) generated the score sequences and their frequencies up to $n=9$. Zeilberger (2016) extended MacMahon’s work and generated the nondecreasing score sequences and their frequencies for $n\leq 15$ using the Maple program. In a survey paper, Guy (1984) notes that Epstein’s problem on $r_{n}(M_{1})$ remains unsolved. In a recent publication, Malinovsky and Moon (2024) proved (2). They used a method which was introduced by Erdős and Wilson (1977) that considered the analogous problem for the vertices of maximum degree in a random labelled graph in which pairs of distinct vertices are joined by an edge with probability $1/2$. In the next paragraph, we outline this method, which was applied to model $M_{1}.$
Set
$s^{\star}_{n}=\max\left\{s_{1}(n),\ldots,s_{n}(n)\right\},$
and
$W_{n}(t)=\sum_{1\leq v<u\leq n}I(t<s_{u}(n)=s_{v}(n)).$ (3)
Under model $M$, we need to identify the value $t_{n}$, if such a value exists, so that
$\lim_{n\rightarrow\infty}\mathsf{P}\left(s^{\star}_{n}>t_{n}\right)=1$ (4)
and
$\lim_{n\rightarrow\infty}\mathsf{P}\left(W_{n}(t_{n})=0\right)=1.$} (5)
Combining, (4) and (5), we obtain
$\lim_{n\rightarrow\infty}r_{n}(M)=1.$} (6)
The first question that could be asked is whether a similar result holds for a more general model, e.g. a chess tournament model with equally strong players, where a player scores one point for a game won and half a point for a game drawn (hereafter model $M_{2}$.) MacMahon (1923b) considered such a model motivated by the Master’s Chess Tournament which was held in London in 1922. In that work, MacMahon generated the nondecreasing score sequences and their frequencies for $n\leq 6$. It is interesting to note that Zermelo (1929) proposed a maximum likelihood estimation of a player’s strength in chess tournaments (see also David and Edwards (2001), pages 161-186.) Zermelo applied his estimation technique in analyzing the 1924 New York Chess Tournament, a famous tournament featuring prominent players like José Raúl Capablanca, Emanuel Lasker, and Alexander Alekhine.
In fact, we can consider an even more general model defined below.
Model $M_{k}$: For integer $k\geq 1$, let $D_{k}=\left\{0,\frac{1}{k},\frac{2}{k},\ldots,\frac{k-2}{k},\frac{k-1}{k},1\right\}$ be the set of all possible values of $X_{ij}$. Each of these values can be obtained with positive probability. For $i\neq j$, $X_{ij}+X_{ji}=1$, $X_{ij}\in D_{k}$. This together with the assumption that all players are equally strong (i.e., $X_{ij}$ are equally distributed), is equivalent to
$0<\mathsf{P}\left(X_{ij}=a\right)=\mathsf{P}\left(X_{ij}=1-a\right)\,\,\,\,\text{for any}\,\,\,\,a\in D_{k}.$ (7)
In this paper, we show that in the round-robin tournament model with equally strong players, where scores take values on a lattice (model $M_{k}$, $k\geq 1$), the probability that there is a unique vertex with the maximum score tends to $1$ as $n$ tends to infinity. In the following section we present some preliminary results. Afterwards, we present a proof of the main result. Finally, the discussion section presents related open problem.
2 Preliminary Results
Let
$\mu=E(X_{ij}),\,\,\,\sigma=(Var(X_{ij}))^{1/2}.$ (8)
Next, let
$t_{n}=(n-1)\mu+x_{n}(n-1)^{1/2}\sigma,$ (9)
where
$x_{n}=[2\log(n-1)-(1+\epsilon)\log(\log(n-1))]^{1/2},\,\,\,\epsilon>0.$ (10)
Assuming (7), we obtain $\mu=\tfrac{1}{2}$ and $\sigma\leq\tfrac{1}{2}$.
Proposition 1.
Assuming Model $M_{k}$, we have
$\lim_{n\rightarrow\infty}\mathsf{P}\left(s^{\star}_{n}>t_{n}\right)=1.$ (11)
Proof.
Since $x_{n}=o\left((n-1)^{1/6}\right)$}, $\frac{x_{n}^{3}}{(n-1)^{1/2}}\rightarrow 0$} and $\sigma(n-1)^{1/2}\rightarrow\infty$, as $n\rightarrow\infty$, it follows from (Feller, 1971, p. 553, Theorem 3) that
$\mathsf{P}(s_{1}(n)>t_{n})\sim 1-\Phi(x_{n}),$ (12)
where $\Phi()$ is the standard normal distribution function. Also, since $x_{n}\rightarrow\infty$ as $n\rightarrow\infty$ (see for example, (Feller, 1968, p. 175, Lemma 2)),
$1-\Phi(x_{n})\sim\frac{1}{x_{n}}\varphi(x_{n}),$ (13)
where $\varphi()$ is the PDF of a standard normal random variable. Combining (12) and (13), we obtain
$\mathsf{P}(s_{1}(n)>t_{n})\sim\frac{(\log(n-1))^{\epsilon/2}}{\sqrt{4\pi}(n-1)}.$ (14)
It was shown in Malinovsky and Rinott (2023) that
$\mathsf{P}\left(s_{1}(n)\leq x,s_{2}(n)\leq x,\ldots,s_{n}(n)\leq x\right)\leq\mathsf{P}\left(s_{1}(n)\leq x\right)\mathsf{P}\left(s_{2}(n)\leq x\right)\cdots\mathsf{P}\left(s_{n}(n)\leq x\right).$ (15)
Taking into account (15), the inequality $1-c\leq e^{-c}$, and the fact that the scores are equally distributed, we obtain
$\mathsf{P}\left(s^{\star}_{n}>t_{n}\right)=1-\mathsf{P}\left(s^{\star}_{n}\leq t_{n}\right)\geq 1-\left(1-\mathsf{P}\left(s_{1}(n)>t_{n}\right)\right)^{n}\geq 1-e^{-n\mathsf{P}\left(s_{1}(n)>t_{n}\right)}.$ (16)
Finally, combining (14) and (16), we obtain
$\lim_{n\rightarrow\infty}\mathsf{P}\left(s^{\star}_{n}>t_{n}\right)=1.$ (17)
Recall the definition of $t_{n}$ in (9) and define
$A_{n}=\left\{\frac{j}{k}:t_{n}-1<\frac{j}{k}\leq n-2\quad\text{and}\quad\frac{j}{k}\text{ is an attainable value of }s_{u}(n-1)\right\},$ (18)
where $s_{u}(n-1)\in\left\{0,\frac{1}{k},\frac{2}{k},\ldots,\frac{(n-2)k}{k}=n-2\right\}$.
Recalling the definition of $W_{n}(t)$ in (3), we have the following proposition.
Proposition 2.
Assuming Model $M_{k}$ and $n\geq 3$, we have
${\mathsf{E}}\left(W_{n}(t_{n})\right)\leq\frac{n(n-1)}{2}\sum_{h\in A_{n}}\mathsf{P}\left(s_{u}(n-1)=h\right)\mathsf{P}\left(s_{v}(n-1)=h\right),$
where the set $A_{n}$ is defined in (18).
Proof.
From the fact that scores are equally distributed and the law of total probability, we obtain
${\mathsf{E}}\left(W_{n}(t_{n})\right)=\frac{n(n-1)}{2}\mathsf{P}\left(t_{n}<s_{u}(n)=s_{v}(n)\right)$
$=\frac{n(n-1)}{2}\sum_{{h\in\left\{\frac{j}{k}:\,\,t_{n}<\frac{j}{k}\leq n-1\right\}}}\mathsf{P}\left(s_{u}(n)=h,s_{v}(n)=h\right)$
$=\frac{n(n-1)}{2}\sum_{a\in D_{k}}\sum_{{{h\in\left\{a+\frac{j}{k}:\,\,t_{n}<a+\frac{j}{k}\leq a+n-2\right\}}}}\mathsf{P}\left(s_{u}(n)=h,s_{v}(n)=h\,|\,X_{uv}=a\right)\mathsf{P}\left(X_{uv}=a\right)$
$=\frac{n(n-1)}{2}\sum_{a\in D_{k}}\sum_{{{h\in\left\{a+\frac{j}{k}:\,\,t_{n}<a+\frac{j}{k}\leq a+n-2\right\}}}}\mathsf{P}\left(s_{u}(n-1)=h-a\right)\mathsf{P}\left(s_{v}(n-1)=h-(1-a)\right)\mathsf{P}\left(X_{uv}=a\right)$
$=\frac{n(n-1)}{2}\sum_{a\in D_{k}}\mathsf{P}\left(X_{uv}=a\right)\sum_{{{h\in\left\{a+\frac{j}{k}:\,\,t_{n}<a+\frac{j}{k}\leq a+n-2\right\}}}}\mathsf{P}\left(s_{u}(n-1)=h-a\right)\mathsf{P}\left(s_{v}(n-1)=h-(1-a)\right)$
$\leq\frac{n(n-1)}{2}\sum_{a\in D_{k}}\mathsf{P}\left(X_{uv}=a\right)\sum_{{{h\in\left\{a+\frac{j}{k}:\,\,t_{n}<a+\frac{j}{k}\leq a+n-2\right\}}}}\mathsf{P}\left(s_{u}(n-1)=h-a\right)\mathsf{P}\left(s_{v}(n-1)=h-a\right)$
$\leq\frac{n(n-1)}{2}\sum_{a\in D_{k}}\mathsf{P}\left(X_{uv}=a\right)\sum_{{{h\in\left\{\frac{j}{k}:\,\,t_{n}-1<\frac{j}{k}\leq n-2\right\}}}}\mathsf{P}\left(s_{u}(n-1)=h\right)\mathsf{P}\left(s_{v}(n-1)=h\right)$
$=\frac{n(n-1)}{2}\sum_{h\in A_{n}}\mathsf{P}\left(s_{u}(n-1)=h\right)\mathsf{P}\left(s_{v}(n-1)=h\right),$ (19)
where the first inequality follows from combining assumption (7) with the rearrangement inequality, and the second inequality follows from adding nonnegative summands. ∎
Proposition 3.
Assuming Model $M_{k}$, we have
$\lim_{n\rightarrow\infty}\mathsf{P}\left(W_{n}(t_{n})=0\right)=1.$} (20)
Proof.
Recall that $X_{ij}\in D_{k}=\left\{0,\frac{1}{k},\frac{2}{k},\ldots,\frac{k-2}{k},\frac{k-1}{k},1\right\}$. Set
$Y_{ij}=kX_{ij}\in\left\{0,1,2,\ldots,k\right\},\,\,\,i\neq j\in\left\{1,2,\ldots,n\right\}.$ (21)
Therefore, recalling notation (1), we have
$ks_{i}(n-1)=\sum_{j=1,j\neq i}^{n-1}{Y_{ij}},i=1,\ldots,n-1.$ (22)
Define $y$ by
$(n-2)\mu+y(n-2)^{1/2}\sigma=\text{an attainable value of }s_{u}(n-1),$ (23)
and for every $y$ satisfying condition (23), define
$P_{n}(y)=\mathsf{P}\Big(s_{u}(n-1)=(n-2)\mu+y(n-2)^{1/2}\sigma\Big).$ (24)
We need the theorem due to Esseen (1945), which we present below, to prove the subsequent Claims 1 and 2.
Theorem 1.
Esseen (1945) (see also (Petrov, 1975, Theorem 6, p. 197 and p. 322)) Let $\left\{X_{n}\right\}$ be a sequence of independent integer-valued random variables having a common distribution with non-zero variance $\sigma^{2}$ and finite third-order moment. If the maximum span of $X_{1}$ is equal 1, then
$\sup_{N}\left|\sigma\sqrt{n}\mathsf{P}\left(\sum_{j=1}^{n}X_{j}=N\right)-\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\left(\frac{N-nE(X_{1})}{\sigma\sqrt{n}}\right)^{2}}\right|=O\left(\frac{1}{\sqrt{n}}\right).$
In the notation of (21) and (22), Theorem 1 asserts that
$\lim_{n\rightarrow\infty}\sup_{N}\left|\sigma\sqrt{n-2}\mathsf{P}(ks_{1}(n-1)=kN)-\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\left(\frac{kN-k(n-2)\mu}{k\sigma\sqrt{n-2}}\right)^{2}}\right|=0.$ (25)
The following two claims are direct consequences of (25).
Claim 1.
For any $y_{1}$ and $y_{2}$ satisfying (23) with $y_{1}\leq y_{2}$, we have $P_{n}(y_{1})\geq P_{n}(y_{2})$ as $n\to\infty$.
Proof.
[Claim 1] From (25) it follows that
$P_{n}(y)\sim\frac{1}{\sigma\sqrt{n-2}}\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}y^{2}}.$ (26)
Claim 1 then follows from (26). ∎
Denote by $y^{\star}$ the smallest element of the set $A_{n}$. It follows from (9) and the subsequent discussion
$y^{\star}=\frac{\lceil(n-1)\frac{k}{2}+x_{n}(n-1)^{1/2}c_{k}-k\rceil}{k},$ (27)
for some $c_{k}\leq\frac{k}{2}$. Appealing again to (25) we obtain the following claim.
Claim 2.
For $y^{\star}$ defined in (27),
$\mathsf{P}\left(ks_{1}(n-1)=ky^{\star}\right)\sim\frac{1}{\sigma\sqrt{n-2}}\frac{(\log(n-1))^{(1+\epsilon)/2}}{\sqrt{2\pi}(n-1)},$ (28)
as $n\rightarrow\infty.$
Now, by Proposition 2, together with Claims 1 and 2, it follows that, as $n\to\infty,$
${\mathsf{E}}\left(W_{n}(t_{n})\right)\leq\frac{n(n-1)}{2}\sum_{h\in A_{n}}\mathsf{P}\left(s_{u}(n-1)=h\right)\mathsf{P}\left(s_{v}(n-1)=h\right)$
$\leq\frac{n(n-1)}{2}\mathsf{P}\left(s_{u}(n-1)=h^{\star}\right)\sum_{h\in A_{n}}\mathsf{P}\left(s_{v}(n-1)=h\right)$
$=\frac{n(n-1)}{2}\mathsf{P}\left(s_{u}(n-1)=h^{\star}\right)\mathsf{P}\left(s_{v}(n-1)>t_{n}-1\right)$
$\sim\frac{n(n-1)}{2}\frac{1}{\sigma\sqrt{n-2}}\frac{(\log(n-1))^{(1+\epsilon)/2}}{\sqrt{2\pi}(n-1)}\mathsf{P}\left(s_{v}(n-1)>t_{n}-1\right)$
$\sim\frac{n(n-1)}{2}\frac{1}{\sigma\sqrt{n-2}}\frac{(\log(n-1))^{(1+\epsilon)/2}}{\sqrt{2\pi}(n-1)}\frac{(\log(n-1))^{\epsilon/2}}{\sqrt{4\pi}(n-1)},$ (29)
where the last line is obtained in a manner similar to (14).
Since $W_{n}(t_{n})\geq 0$, and by (2), it follows that
$\lim_{n\rightarrow\infty}{\mathsf{E}}\left(W_{n}(t_{n})\right)=0.$} (30)
3 Main Result
Result 1.
For any fixed $k\geq 1$,
$\lim_{n\rightarrow\infty}r_{n}(M_{k})=1.$} (31)
Proof.
Appealing to Proposition 3 and using the inequality $W_{n}=W_{n}I(W_{n}>0)\geq I(W_{n}>0)\geq 0$, we obtain (5) under Model $M_{k}$, i.e.,
$\lim_{n\to\infty}\mathsf{P}\bigl(W_{n}(t_{n})=0\bigr)=1.$ (32)
Combining Propositions 1 and (32), we obtain Result 1. ∎
4 Discussion
A possible extension of the model $M_{k}$ with the finite number of payoffs is the extension to the countable number of payoffs in the interval $[0,1]$. Such model is formally described as follows. Let $D$ in $[0,1]$ be the discrete set of all possible values of $X_{ij}$. Each of these values can be obtained with positive probability. For $i\neq j$, $X_{ij}+X_{ji}=1$, $X_{ij}\in D$. This together with the assumption that all players are equally strong (i.e., $X_{ij}$ are equally distributed) is equivalent to $0<\mathsf{P}\left(X_{ij}=a\right)=\mathsf{P}\left(X_{ij}=1-a\right)\,\,\,\,\text{for any}\,\,\,\,a\in D.$ Under such a model, Propositions 1 and 2 still hold. However, the local limit theorem used in the proof of Proposition 3 (Claims 1 and 2), which holds for integer-valued random variables, should be extended to this case. One possible direction is to derive the local limit theorems, as was done in number theory (Alon and Freiman, 1988), by using the circle method.
Acknowledgements
I would like to thank Noga Alon for the discussions and suggestions. I would also like to thank Doron Zeilberger for the opportunity to present a talk at his seminar and for his interesting questions.